# Extreme Miscalibration and the Illusion of Adversarial Robustness

### v1

@@ F.4 Transferability of High Temperature Training
It is shown that training with a high temperature leads to a consistent gain in adversarial robustness to unseen adversarial attack forms. However, an adversary may attempt to exploit attack transferability when looking to attack the target model trained with high temperature. To explore this notion of a transfer attack, with the rt dataset, Table 13 shows the impact of finding adversarial examples for the source baseline $*{}^{*}$ model and assessing their efficacy on the target baseline $*{}^{*}$ $\oplus\ T$ model. It is evident from the significant increase in the adversarial accuracy for all the attack forms (bae, tf, pwws, and dg), that a transfer attack is not able to degrade the observed robustness gains for models trained with high temperatures.
@@
$*{}^{*}$ $*{}^{*}$ $\underset{\pm 1.20}{31.39}$ $\underset{\pm 0.49}{17.82}$ $\underset{\pm 0.62}{20.42}$ $\underset{\pm 0.94}{20.11}$
@@
$*{}^{*}$ $\oplus\ T$ $\underset{\pm 0.30}{50.13}$ $\underset{\pm 0.38}{46.90}$ $\underset{\pm 1.22}{47.53}$ $\underset{\pm 1.13}{46.09}$
Table 22: Transferability: adversarial examples are found for the source model and evaluated on the target model on the rt test set. The results here demonstrate that the standard trained, high temperature ($\oplus T$) model’s robustness gains relative to the $*{}^{*}$ model cannot be compromised by a transferability attack, i.e. the performance of the $\oplus T$ model are not degraded by adversarial examples generated from the $*{}^{*}$ model. Test-time calibration (cal) is used to mitigate IOR.
@@ F.5 Detailed Performance Breakdown
Figure 4 presents the adversarial accuracy of baseline $*{}^{*}$ models trained with different training temperatures. In this section, for reference, we provide the detailed breakdown (average across 3 seeds and standard deviation) of performances for the different training temperatures for each dataset: rt (Table 23), emotion (Table 24), cola (Table 25), qnli (Table 26), and mrpc (Table 27). These results are given for the DeBERTa model as in the main paper.
@@
Table 23: rt: The use of a training temperature, $T$, is a simple adjustment in standard baseline model training ($*{}^{*}$), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 24: emotion: The use of a training temperature, $T$, is a simple adjustment in standard baseline model training ($*{}^{*}$), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 25: cola: The use of a training temperature, $T$, is a simple adjustment in standard model training ($*{}^{*}$), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 26: qnli: The use of a training temperature, $T$, is a simple adjustment in standard model training ($*{}^{*}$), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 27: mrpc: The use of a training temperature, $T$, is a simple adjustment in standard model training ($*{}^{*}$), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 28: agnews: The use of a training temperature, $T$, is a simple adjustment in standard model training ($*{}^{*}$), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 29: agnews: The use of a training temperature, $T$, is a simple adjustment in standard model training ($*{}^{*}$), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Training with 10k samples - 1/10th of default agnews training set size. Test-time calibration (cal) is used to mitigate IOR.
@@ F.6 Reproducing with Other Models
Table 30: BERT: The use of a training temperature, $T$, is a simple adjustment in standard model training ($*{}^{*}$), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR. Result for rt dataset.
@@
Table 31: RoBERTa: The use of a training temperature, $T$, is a simple adjustment in standard model training ($*{}^{*}$), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR. Result for rt dataset.
@@ F.7 High Temperature Training GradNorm Ablation
$*{}^{*}$ $\oplus T$ $\underset{\pm 0.44}{87.55}$ $\underset{\pm 0.84}{35.83}$ $\underset{\pm 4.57}{26.83}$ $\underset{\pm 3.07}{31.49}$ $\underset{\pm 4.71}{35.18}$
Table 32: Gradnorm ($*{}^{*}$) Ablation. Results are presented for the baseline system for the rt dataset and the DeBERTa model. A high training temperature of $T=200$ is used. Test-time calibration (cal) is used to mitigate IOR.
@@ F.9 Other Ablations
Augmentation based adversarial training approaches, such as $*{}^{*}$ in the main paper, have twice as many training steps (due to there being double the training set size). To match the standard training setting, in Table 34 we evaluated the training with high temperature approach combined with $*{}^{*}$ at half the number of training steps. Similarly, in Table 35 we consider the inverse setting, where we double the number of training iterations for the $*{}^{*}$ model (in standard training), as well as linearly scaling the learning rate scheduler across the increased number of iterations.
@@
$*{}^{*}$ $\oplus T$ default $\underset{\pm 0.44}{87.55}$ $\underset{\pm 3.07}{31.49}$ $\underset{\pm 4.71}{35.18}$
$*{}^{*}$ $\oplus T$ default $\underset{\pm 0.22}{87.09}$ $\underset{\pm 1.67}{31.43}$ $\underset{\pm 1.90}{36.40}$
@@
Table 34: Matched number of iterations for $*{}^{*}$ and high temperature training with $*{}^{*}$ by halving the number of training steps for $*{}^{*}$. Test-time calibration (cal) is used to mitigate IOR.
@@
$*{}^{*}$ 5 $\underset{\pm 0.19}{88.56}$ $\underset{\pm 0.14}{32.40}$ $\underset{\pm 0.48}{18.79}$ $\underset{\pm 1.22}{21.36}$ $\underset{\pm 0.66}{21.11}$
@@
Table 35: Doubling training iterations for the $*{}^{*}$ model with scaled scheduler decay to match number of iterations in augmentation based AT. Test-time calibration (cal) is used to mitigate IOR.

### v3

@@ F.4 Transferability of High Temperature Training
It is shown that training with a high temperature leads to a consistent gain in adversarial robustness to unseen adversarial attack forms. However, an adversary may attempt to exploit attack transferability when looking to attack the target model trained with high temperature. To explore this notion of a transfer attack, with the rt dataset, Table 15 shows the impact of finding adversarial examples for the source baseline ST model and assessing their efficacy on the target baseline ST $\oplus\ T$ model. It is evident from the significant increase in the adversarial accuracy for all the attack forms (bae, tf, pwws, and dg), that a transfer attack is not able to degrade the observed robustness gains for models trained with high temperatures.
@@
baseline baseline $\underset{\pm 1.20}{31.39}$ $\underset{\pm 0.49}{17.82}$ $\underset{\pm 0.62}{20.42}$ $\underset{\pm 0.94}{20.11}$
@@
baseline $\oplus\ T$ $\underset{\pm 0.30}{50.13}$ $\underset{\pm 0.38}{46.90}$ $\underset{\pm 1.22}{47.53}$ $\underset{\pm 1.13}{46.09}$
Table 24: Transferability: adversarial examples are found for the source model and evaluated on the target model on the rt test set. The results here demonstrate that the standard trained, high temperature ($\oplus T$) model’s robustness gains relative to the baseline model cannot be compromised by a transferability attack, i.e. the performance of the $\oplus T$ model are not degraded by adversarial examples generated from the baseline model. Test-time calibration (cal) is used to mitigate IOR.
@@ F.5 Detailed Performance Breakdown
Figure 4 presents the adversarial accuracy of baseline ST models trained with different training temperatures. In this section, for reference, we provide the detailed breakdown (average across 3 seeds and standard deviation) of performances for the different training temperatures for each dataset: rt (Table 25), emotion (Table 26), cola (Table 27), qnli (Table 28), and mrpc (Table 29). These results are given for the DeBERTa model as in the main paper.
@@
Table 25: rt: The use of a training temperature, $T$, is a simple adjustment in standard baseline model training (baseline), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 26: emotion: The use of a training temperature, $T$, is a simple adjustment in standard baseline model training (baseline), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 27: cola: The use of a training temperature, $T$, is a simple adjustment in standard model training (baseline), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 28: qnli: The use of a training temperature, $T$, is a simple adjustment in standard model training (baseline), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 29: mrpc: The use of a training temperature, $T$, is a simple adjustment in standard model training (baseline), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 30: agnews: The use of a training temperature, $T$, is a simple adjustment in standard model training (baseline), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR.
@@
Table 31: agnews: The use of a training temperature, $T$, is a simple adjustment in standard model training (ST), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Training with 10k samples - 1/10th of default agnews training set size. Test-time calibration (cal) is used to mitigate IOR.
@@ F.6 Reproducing with Other Models
Table 32: BERT: The use of a training temperature, $T$, is a simple adjustment in standard model training (ST), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR. Result for rt dataset.
@@
Table 33: RoBERTa: The use of a training temperature, $T$, is a simple adjustment in standard model training (ST), where the temperature parameter, $T$, is used to scale down predicted model logits. Higher training temperatures enhance model robustness against unseen adversarial attacks (bae, tf, pwws, dg) without requiring prior knowledge of these attack forms during training. Results here report the clean and adversarial accuracy. Test-time calibration (cal) is used to mitigate IOR. Result for rt dataset.
@@ F.7 High Temperature Training GradNorm Ablation
baseline $\oplus T$ $\underset{\pm 0.44}{87.55}$ $\underset{\pm 0.84}{35.83}$ $\underset{\pm 4.57}{26.83}$ $\underset{\pm 3.07}{31.49}$ $\underset{\pm 4.71}{35.18}$
Table 34: Gradnorm () Ablation. Results are presented for the baseline system for the rt dataset and the DeBERTa model. A high training temperature of $T=200$ is used. Test-time calibration (cal) is used to mitigate IOR.
@@ F.9 Other Ablations
Augmentation based adversarial training approaches, such as dg-aug in the main paper, have twice as many training steps (due to there being double the training set size). To match the standard training setting, in Table 36 we evaluated the training with high temperature approach combined with dg-aug at half the number of training steps. Similarly, in Table 37 we consider the inverse setting, where we double the number of training iterations for the baseline model (in standard training), as well as linearly scaling the learning rate scheduler across the increased number of iterations.
@@
baseline $\oplus T$ default $\underset{\pm 0.44}{87.55}$ $\underset{\pm 3.07}{31.49}$ $\underset{\pm 4.71}{35.18}$
dg-aug $\oplus T$ default $\underset{\pm 0.22}{87.09}$ $\underset{\pm 1.67}{31.43}$ $\underset{\pm 1.90}{36.40}$
@@
Table 36: Matched number of iterations for baseline and high temperature training with dg-aug by halving the number of training steps for dg-aug. Test-time calibration (cal) is used to mitigate IOR.
@@
baseline 5 $\underset{\pm 0.19}{88.56}$ $\underset{\pm 0.14}{32.40}$ $\underset{\pm 0.48}{18.79}$ $\underset{\pm 1.22}{21.36}$ $\underset{\pm 0.66}{21.11}$
@@
Table 37: Doubling training iterations for the baseline model with scaled scheduler decay to match number of iterations in augmentation based AT. Test-time calibration (cal) is used to mitigate IOR.